OpenStudy (anonymous):

An engineer wants to make 60% cement mixture. How many pounds of a 80% cement mixture must she mix with 20 pounds of a 40% mixture to create a 60% cement mixture?

OpenStudy (anonymous):

you are full of this mix problem, hug? must be algebra word problem hell week

OpenStudy (anonymous):

yes it sucks!

OpenStudy (anonymous):

you want reasoning or equation ?

OpenStudy (anonymous):

idk whatever solves it

OpenStudy (anonymous):

depends on how much you need to understand if we call \(x\) the amount of \(80\%\) mix then you must solve \[.8x+8=.6(20+x)\]

OpenStudy (anonymous):

which is easier to solve if you multiply by ten and write \[8x+80=6(20+x)\]

OpenStudy (anonymous):

20? @satellite73

OpenStudy (anonymous):

idk i didn't do it hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so 20 pounds?

OpenStudy (anonymous):

oui